# Differentially Private Adaptation of Diffusion Models via Noisy Aggregated Embeddings

### v3

@@ Abstract
Figure 2: Overview of DPAgg-TI. We first apply Textual Inversion to extract embeddings for each image in the private dataset. These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{*}_{\text{DP}}$. Finally, images are generated using the corresponding token $S^{*}$>.
1 Introduction
@@
The rapid adoption of diffusion models  (Ho, Jain, and Abbeel 2020; Song et al. 2021; Rombach et al. 2022) has raised significant privacy and legal concerns. These models are vulnerable to privacy attacks, such as membership inference (Duan et al. 2023), where attackers determine if a specific data point was used for training, and data extraction (Carlini et al. 2023), which enables reconstruction of training data. This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact. Additionally, reliance on large datasets scraped without consent raises copyright concerns (Vyas, Kakade, and Barak 2023), as diffusion models can reproduce original artworks without credit or compensation. These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
@@
Differential privacy (DP) (Dwork 2006) is a widely adopted framework for addressing these challenges. One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) (Abadi et al. 2016), which modifies traditional SGD by adding noise to clipped gradients. However, applying DP-SGD to train diffusion models poses several challenges. It introduces significant computational and memory overhead due to per-sample gradient clipping (Hoory et al. 2021), which is essential for bounding gradient sensitivity (Dwork et al. 2006; Abadi et al. 2016). DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis. Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm. Consequently, existing diffusion models trained with DP-SGD are limited to relatively small-scale images. (Dockhorn et al. 2023; Ghalebikesabi et al. 2023).
@@
Independent of privacy concerns, Textual Inversion (TI) (Gal et al. 2023) provides an effective method for adapting diffusion models to specific styles or content without modifying the model. Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs. A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying the application of privacy-preserving mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model. Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
@@
In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD. Standard TI does not offer formal privacy guarantees, so to address this limitation, we introduce a private variant of TI, called Differentially Private Aggregation via Textual Inversion (DPAgg-TI) and summarize it in Figure 2. Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid. This approach ensures efficient and secure adaptation to private datasets.
@@
Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints. Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms (Paris 2024 2023), we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy. We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise. Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality. This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
@@
2 Background and Related Work
2.1 Diffusion Models
Diffusion models (Ho, Jain, and Abbeel 2020; Song et al. 2021; Song, Meng, and Ermon 2021; Rombach et al. 2022) leverage an iterative denoising process to generate high-quality images that align with a given conditional input from random noise. In text-to-image generation, this conditional input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditional format, it is first broken down into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP (Radford et al. 2021), outputting a single text-conditional vector $y$ that serves as the conditioning input. This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
The image generation process, also known as the reverse diffusion process, comprises of $T$ discrete timesteps and starts with pure Gaussian noise $x_{T}$. At each decreasing timestep $t$, the denoising model, which often utilizes a U-Net structure with cross-attention layers, takes a noisy image $x_{t}$ and text conditioning $y$ as inputs and predicts the noise component $\epsilon_{\theta}(x_{t},y,t)$, where $\theta$ denotes the denoising model’s parameters. The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ that aligns with the text conditional $y$.
The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error (MSE) between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$. The denoising model is therefore trained over the following optimization problem:
@@ 2.1 Diffusion Models
Textual Inversion.
Textual Inversion (TI) (Gal et al. 2023) is an adaptation technique that enables personalization using a small dataset of typically 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$. During training, images are conditioned on phrases such as “A photo of SS” or “A painting in the style of SS”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss given in (1) with respect to $u$, while keeping the diffusion model parameters $\theta$ fixed, iteratively refining this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u^{*}$ is formalized as follows:
@@
2.2 Differential Privacy.
In this work, we adopt differential privacy (DP) (Dwork et al. 2006; Dwork 2006) as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
@@
To achieve DP, the Gaussian mechanism is often applied (Dwork, Roth et al. 2014; Balle and Wang 2018), adding Gaussian noise scaled by the sensitivity of the function $f$ and privacy parameters $\varepsilon$ and $\delta$. Specifically, noise with standard deviation $\sigma=\frac{\Delta_{f}\sqrt{2\ln(1.25/\delta)}}{\varepsilon}$ is added to the output [^1] In practice, we use numerical privacy accountant such as (Balle and Wang 2018; Mironov 2017) to calibrate the noise. (Balle and Wang 2018), where $\Delta_{f}$ represents $\ell_{2}$-sensitivity of the target function $f(\cdot)$. When the context is clear, we may omit the subscript $f$. This mechanism enables a smooth privacy-utility tradeoff and is widely used in privacy-preserving machine learning, including in DP-SGD (Abadi et al. 2016), which applies Gaussian noise during model updates to achieve DP.
Privacy Amplification by Subsampling.
Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (such as the Gaussian mechanism) is applied. Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
@@
This result is well-known (Steinke (2022, Theorem 29)), with tighter amplification bounds available for Gaussian mechanisms (Mironov 2017).
2.3 Private Adaptation of Diffusion Models
Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Dockhorn et al. Dockhorn et al. (2023) proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD. Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset, as explored by Ghalebikesabi et al. (2023). While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase (Tramèr, Kamath, and Carlini 2024).
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in the VIP approach (Yu et al. 2024), which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations (Sander et al. 2024), aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
@@ 2.3 Private Adaptation of Diffusion Models
3 Differentially Private Adaptation via Textual Inversion
Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards. Instead of training a single token embedding on the entire dataset as in regular TI, we train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$, as illustrated in Figure 2. We can formalize the encoding process as follows:
Then, we can aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating the centroid. The purpose of this aggregation is to limit the sensitivity of the final output to each $x^{(i)}$. In order to provide DP guarantees, we also add isotropic Gaussian noise to the centroid. We can therefore define the resulting embedding vector $u^{*}_{\text{DP}}$ as follows:
$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (5)
where the minimum $\sigma$ required to provide $(\varepsilon,\delta)$-DP is given by the following expression based on Balle and Wang (2018, Theorem 1):
$\sigma=\frac{\Delta}{n}\cdot\frac{\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (6)
In the context of our problem, $\Delta=\sup_{i,j}\|u^{(i)}-u^{(j)}\|$. Since our embedding vectors are directional, we can normalize each $u^{(i)}$, allowing us to set $\Delta=2$.
To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors. Then the standard privacy amplification by subsampling bounds (such as (3)) can be applied. Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute the output embedding as follows:
$u^{*}_{\text{DP}}=\frac{1}{m}\sum_{u^{(i)}\in D_{\text{sub}}}u^{(i)}+\mathcal{N}(0,\sigma^{2}I),$ (7)
where $\sigma$ can be computed numerically for any target $\varepsilon,\delta$ and subsampling rate $\frac{m}{n}$.
4 Experimental Results
4.1 Datasets
The second dataset contains 47 pictograms from the Paris 2024 Olympics (Paris 2024 2023), permitted strictly for non-commercial editorial use (International Olympic Committee 2025). These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data. This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
4.2 Style Transfer Results
@@
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI (Gal et al. 2023) embeddings on Stable Diffusion v1.5 (Rombach et al. 2022) using DPAgg-TI. Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience. For regular TI, we utilize the default process to embed the private dataset without any additional noise. For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
@@
Figures 4 and  5 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$. We used the same random seed to generate embeddings, subsample images, and sample DP noise for ease of visual comparison between different configurations. As with common practice, we set $\delta=1/n$. Since $\sigma$ is undefined for $\varepsilon=0$, we demonstrate the results of $\varepsilon\approx 0$, in other words, infinite noise, by setting $\varepsilon=10^{-5}$. The purpose of this parameter value is to demonstrate the image generated when $u^{*}_{\text{DP}}$ contains zero information about the target dataset. Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work, while those of Paris 2024 pictograms captured the logo’s original structure. In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in diminished stylistic fidelity.
@@
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}_{\text{DP}}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}_{\text{DP}}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements, or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms). With this in mind, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards infinity, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees. We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 4 and intermediate $\varepsilon$ values in Figure 5) which restores as $u^{*}_{\text{DP}}$ drifts even further from its optimal value. We hypothesize that this is due to drifted $u^{*}_{\text{DP}}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}_{\text{DP}}$ to be disregarded from $y_{u^{*}_{\text{DP}}}$ and the prompt fidelity to be restored. Another possible explanation is that the temporary drop in prompt fidelity is due to the drift path of $u^{*}_{\text{DP}}$ passing through non-linear regions within embedding space. We leave further investigations into this observation for future work.

### v4

@@ Abstract
Figure 2: Overview of DPAgg-TI. We first apply Textual Inversion to extract embeddings for each image in the private dataset. These embeddings are then aggregated with differentially private mechansim, incorporating subsampling to produce a private embedding $u^{*}_{\text{DP}}$. Finally, images are generated using the corresponding token $S^{*}$.
I Introduction
@@
The rapid adoption of diffusion models  [1, 2, 3] has raised significant privacy and legal concerns. These models are vulnerable to privacy attacks, such as membership inference [4], where attackers determine if a specific data point was used for training, and data extraction [5], which enables reconstruction of training data. This risk is amplified during fine-tuning on smaller, domain-specific datasets, where each record has a greater impact. Additionally, reliance on large datasets scraped without consent raises copyright concerns [6], as diffusion models can reproduce original artworks without credit or compensation. These issues highlight the urgent need for privacy-preserving technologies and clearer ethical and legal guidelines for generative models.
@@
Differential privacy (DP) [7] is a widely adopted framework for addressing these challenges. One standard approach for ensuring DP in deep learning is Differentially Private Stochastic Gradient Descent (DP-SGD) [8], which modifies traditional SGD by adding noise to clipped gradients. However, applying DP-SGD to train diffusion models poses several challenges. It introduces significant computational and memory overhead due to per-sample gradient clipping [9], which is essential for bounding gradient sensitivity [10, 8]. DP-SGD is also incompatible with batch-wise operations like batch normalization, as these link samples and hinder sensitivity analysis. Furthermore, training large models with DP-SGD often leads to substantial performance degradation, particularly under realistic privacy budgets since the required noise scales with the gradient norm. Consequently, existing diffusion models trained with DP-SGD are limited to small-scale images [11, 12].
@@
Independent of privacy concerns, Textual Inversion (TI) [13] effectively adapts diffusion models to specific styles or content without modifying the model. Instead, TI learns an external embedding vector that captures the style or content of a target image set, which is then incorporated into text prompts to guide the model’s outputs. A key advantage of TI is its ability to compress a style into a compact vector, reducing computational and memory demands while simplifying privacy mechanisms, as privacy constraints can be applied directly to embeddings rather than the full model. Additionally, since TI avoids direct model optimization, it remains efficient and compatible with DP constraints on smaller datasets.
@@
In this work, we propose a novel privacy-preserving adaptation method for smaller datasets, leveraging TI to avoid the extensive model updates required by DP-SGD. Standard TI does not offer formal privacy guarantees, so we introduce a private variant, Differentially Private Aggregation via Textual Inversion (DPAgg-TI), summarized in Figure 2. Our method decouples interactions among samples by learning a separate embedding for each target image, which are then aggregated into a noisy centroid. This approach ensures efficient and secure adaptation to private datasets.
@@
Our experiments demonstrate the effectiveness of DPAgg-TI, showing that TI remains robust in preserving stylistic fidelity even under privacy constraints (Figure 1). Applying our method to a private artwork collection by @eveismyname and Paris 2024 Olympics pictograms [14], we show that DPAgg-TI captures nuanced stylistic elements while ensuring privacy. We observe a trade-off between privacy (controlled by DP parameter $\varepsilon$) and image quality: lower $\varepsilon$ reduces fidelity but maintains the target style under moderate noise. Subsampling further amplifies privacy by reducing sensitivity to individual data points, mitigating noise impact on image quality. This framework enables privacy-preserving adaptation of diffusion models to new styles and domains while protecting sensitive data.
@@
II Background and Related Work
II-A Diffusion Models
Diffusion models [1, 2, 15, 3] leverage an iterative denoising process to generate high-quality images that align with a given conditioning input from random noise. In text-to-image generation, this conditioning input is based on a textual description (a prompt) that guides the model in shaping the image to reflect the content and style specified by the text. To convert the text prompt into a suitable conditioning format, it is first broken down into discrete tokens, each representing a word or sub-word unit. These tokens are then converted into a sequence of embedding vectors $v_{i}$ that encapsulate the meaning of each token within the model’s semantic space. Next, these embeddings pass through a transformer text encoder, such as CLIP [16], outputting a single text-conditioning vector $y$ that serves as the conditioning input. This vector $y$ is then incorporated at each denoising step, guiding the model to align the output image with the specific details outlined in the prompt.
The image generation process, also known as the reverse diffusion process, comprises of $T$ discrete timesteps and starts with pure Gaussian noise $x_{T}$. At each decreasing timestep $t$, the denoising model, which often utilizes a U-Net structure with cross-attention layers, takes a noisy image $x_{t}$ and text conditioning $y$ as inputs and predicts the noise component $\epsilon_{\theta}(x_{t},y,t)$, where $\theta$ denotes the denoising model’s parameters. The predicted noise is then used to make a reverse diffusion step from $x_{t}$ to $x_{t-1}$, iteratively refining the noisy image closer to a coherent output $x_{0}$ conditioned on $y$.
The objective function for a text-conditioned diffusion model, given both the noisy image $x_{t}$ and the text conditioning $y$, is typically a mean squared error between the true noise $\epsilon$ and the predicted noise $\epsilon_{\theta}(x_{t},y,t)$. The denoising model is therefore trained over the optimization problem
@@ II-A Diffusion Models
II-B Textual Inversion
Textual Inversion (TI) [13] is an adaptation technique that enables personalization using a small dataset of typically 3-5 images. This approach essentially learns a new token that encapsulates the semantic meaning of the training images, allowing the model to associate specific visual features with a custom token.
To achieve this, TI trains a new token embedding, denoted as $u$, representing a placeholder token, denoted as $S$. During training, images are conditioned on phrases such as “A photo of SS” or “A painting in the style of SS”. However, unlike the fixed embeddings of typical tokens $v_{i}$, $u$ is a learnable parameter. Let $y_{u}$ denote the text conditioning vector resulting from a prompt containing the token $S$. Through gradient descent, TI minimizes the diffusion model loss with respect to $u$, instead of the diffusion model parameters $\theta$, which we keep fixed. By doing so, we iteratively refine this embedding to capture the unique characteristics of the training images. The resulting optimal embedding $u^{*}$ is formalized as
@@ II-B Textual Inversion
II-C Differential Privacy
In this work, we adopt differential privacy (DP) [10, 7] as our privacy framework. Over the past decade, DP has become the gold standard for privacy protection in both research and industry. It measures the stability of a randomized algorithm with respect to changes in an input instance, thereby quantifying the extent to which an adversary can infer the existence of a specific input based on the algorithm’s output.
@@ II-C Differential Privacy
To achieve DP, the Gaussian mechanism is commonly applied [17, 18], adding Gaussian noise scaled by the $\ell_{2}$-sensitivity $\Delta$ and privacy parameters $(\varepsilon,\delta)$. We add zero-mean isotropic Gaussian noise with standard deviation
$\sigma\;=\;\frac{\Delta\,\sqrt{2\ln(1.25/\delta)}}{\varepsilon}.$ (3)
In practice, we calibrate $\sigma$ using numerical privacy accountants (e.g., the analytic Gaussian mechanism and RDP) [18, 19]. This mechanism enables a smooth privacy–utility trade-off and is widely used in privacy-preserving machine learning, including DP-SGD [8], which adds Gaussian noise to model updates to achieve DP.
II-C1 Privacy Amplification by Subsampling
Subsampling is a standard technique in DP, where a full dataset of size $n$ is first subsampled to $m$ records without replacement (typically with $m\ll n$) before the privatization mechanism (e.g. the Gaussian mechanism) is applied. Specifically, if a mechanism provides $(\varepsilon,\delta)$-DP on a dataset of size $m$, it achieves $(\varepsilon^{\prime},\delta^{\prime})$-DP on the subsampled dataset, where $\delta^{\prime}=\frac{m}{n}\delta$ and
@@ II-C1 Privacy Amplification by Subsampling
This result is well-known [20, Theorem 29], with tighter amplification bounds available for Gaussian mechanisms [19].
II-D Private Adaptation of Diffusion Models
Recent advancements in applying DP to diffusion models have aimed to balance privacy preservation with the high utility of generative outputs. Early work on differentially private generative models, such as Chen et al.’s [21] investigation of DP-GANs with model inversion defenses, established foundational principles for protecting generative models from privacy breaches during training. Dockhorn et al. [11] proposed a Differentially Private Diffusion Model (DPDM) that enables privacy-preserving generation of realistic samples, setting a foundational approach for adapting diffusion processes using DP-SGD. Another common strategy involves training a model on a large public dataset, followed by differentially private fine-tuning on a private dataset [12]. While effective in certain contexts, this approach raises privacy concerns, particularly around risks of information leakage during the fine-tuning phase [22].
In response to these limitations, various adaptation techniques have emerged. Although not specific to diffusion models, some methods focus on training models on synthetic data followed by DP-constrained fine-tuning, as in Yu et al. [23], which demonstrates the feasibility of applying DP in later adaptation stages. Other approaches explore differentially private learning of feature representations [24], aiming to distill private information into a generalized embedding space while maintaining DP guarantees. Although these adaptations are not yet implemented for diffusion models, they lay essential groundwork for developing secure and efficient privacy-preserving generative models.
@@ II-D Private Adaptation of Diffusion Models
III Differentially Private Adaptation via Textual Inversion
Let $x^{(1)},\dots,x^{(n)}$ represent a target dataset of images whose characteristics we wish to privately adapt our image generation towards. Instead of training a single token embedding on the entire dataset as in regular TI, we train a separate embedding $u^{(i)}$ on each $x^{(i)}$ to obtain a set of embeddings $u^{(1)},\dots,u^{(n)}$, as illustrated in Figure 2. We can formalize the encoding process as
Then, we aggregate the embeddings $u^{(1)},\dots,u^{(n)}$ by calculating their centroid and adding isotropic Gaussian noise. To ensure bounded sensitivity, we employ a purely directional token embedding (semantics depend only on direction), such as CLIP [16], and $\ell_{2}$-normalize each embedding vector prior to aggregation. We can therefore define the resulting centroid $u^{*}_{\text{DP}}$ as
$u^{*}_{\text{DP}}=\frac{1}{n}\sum_{i=1}^{n}\frac{u^{(i)}}{\|u^{(i)}\|}+\mathcal{N}(0,\sigma^{2}I),$ (6)
where $\sigma$ is given by (3).
Under this normalization, the $\ell_{2}$-sensitivity of $u^{*}_{\text{DP}}$ is
$\Delta=\frac{1}{n}\sup_{u,u^{\prime}}\left\|\frac{u}{\|u\|}-\frac{u^{\prime}}{\|u^{\prime}\|}\right\|=\frac{2}{n},$ (7)
To reduce the amount of noise needed to provide the same level of DP, we employ subsampling: instead of computing the centroid over all $n$ embedding vectors, we randomly sample $m\leq n$ embedding vectors without replacement and compute the centroid over only the sampled vectors. Then the standard privacy amplification by subsampling bounds (such as (4)) can be applied. Formally, we sample $D_{\text{sub}}\subseteq\{u^{(1)},\dots,u^{(n)}\}$ where $|D_{\text{sub}}|=m$, and compute $u^{*}_{\text{DP}}$ as
[…]
where $\sigma$ can be computed numerically using (3) with $\varepsilon^{\prime}$ and $\delta^{\prime}$ from Section II-C1.
IV Experimental Results
IV-A Datasets
The second dataset contains 47 pictograms from the Paris 2024 Olympics [14], permitted strictly for non-commercial editorial use [25]. These pictograms were officially released in February 2023, several months after the release of Stable Diffusion v1.5, ensuring they were absent from the model’s pretraining data. This dataset allows us to assess how well our approach adapts to newly introduced visual styles that the base model has never encountered.
@@ IV-A Datasets
IV-B Style Transfer Results
@@
Using both the @eveismyname and Paris 2024 pictograms dataset, we trained TI [13] embeddings on Stable Diffusion v1.5 [3] using DPAgg-TI. Our primary goal is to investigate how DP configurations, specifically the privacy budget $\varepsilon$ and subsampling size $m$, affect the generated images quality and privacy resilience. For regular TI, we utilize the default process to embed the private dataset without any additional noise. For the DPAgg-TI, we test multiple configurations of $m$ and $\varepsilon$ to analyze the trade-off between image fidelity and privacy.
Figures 4 and 5 present generated images across two key configurations: (1) regular TI without DP, (2) DPAgg-TI with DP at different values of $m$ and $\varepsilon$. To ensure reproducibility and fair comparison across all experimental conditions, we fixed the random seed for the entire generation pipeline. This design choice allows us to isolate the effect of our style transfer method while holding other sources of randomness constant. As with common practice, we set $\delta=1/n$.
@@
Images generated without DP closely resemble the unique stylistic elements of the target dataset. In particular, images adapted using @eveismyname images displayed crisp details and nuanced color gradients characteristic of the artist’s work (Figure 4), while those of Paris 2024 pictograms captured the logo’s original structure (Figure 5). In contrast, DP configurations introduce a discernible degradation in image quality, with lower epsilon values and smaller subsampling sizes resulting in diminished stylistic fidelity.
As a no-learning baseline, we consider the limit $\varepsilon\to 0$, under which $u^{*}_{\text{DP}}$ should convey zero information about the target dataset. Since $\sigma\propto 1/\varepsilon$ is undefined at $\varepsilon=0$, we approximate this regime by setting $\varepsilon=10^{-5}\approx 0$, which yields effectively infinite noise.
As $\varepsilon\rightarrow 0$, the resulting token embedding $u^{*}_{\text{DP}}$ gradually loses its semantic meaning, leading to a loss of stylistic fidelity. In particular, $y_{u^{*}_{\text{DP}}}$ tends towards $y$ (a conditioning vector independent of the learnable embedding). In our results, this manifests as a painting of Taylor Swift devoid of the artist-specific stylistic elements (Figure 4), or a generic icon of a dragon (with color, as opposed to the black and white design of the pictograms, Figure 5). To verify this interpretation, we generated images with the same prompts but without the special token $S^{*}$ and compared them to the $\varepsilon\approx 0$ generations. The images were visually identical, confirming that at $\varepsilon\approx 0$, the token becomes semantically meaningless and is ignored by the text encoder.
@@
With this in mind, given a fixed seed, $\varepsilon$ can be interpreted as a drift parameter, representing the progression from the optimal $u^{*}_{\text{DP}}$ towards a semantically meaningless embedding, gradually steering the generated image away from the target style in exchange for stronger privacy guarantees. We also observe instances where there is a temporary drop in prompt fidelity (e.g., $m=16,\varepsilon\in[0.5,1]$ in Figure 4 and intermediate $\varepsilon$ values in Figure 5) which restores as $u^{*}_{\text{DP}}$ drifts even further from its optimal value. We hypothesize that this is due to drifted $u^{*}_{\text{DP}}$ capturing a different meaning unrelated to the prompt, before losing any meaning that could be interpreted by Stable Diffusion’s text encoder, causing $u^{*}_{\text{DP}}$ to be disregarded from $y_{u^{*}_{\text{DP}}}$ and the prompt fidelity to be restored.